List of participants

AGLIARDI Rossela, Dipartimento di Matematica, Via Machiavelli 35, 44100 Ferrara, ITALY

E-mail: agl@ifeuniv.unife.it

ALBINUS Günter, Institut für Angewandte Analysis und Stochastik, Mohrenstr. 39, D-10117 Berlin, GERMANY

E-mail: albinus@iaas-berlin.d400.de

APUSHKINSKAYA Darya, ap. 14, Dalnevostochnyi pr. 80, St. Peterburg 193224, RUSSIA

BARDARO Carlo, Dipartimento di Matematica, Università degli Studi di Perugia, Via Vanvitelli 1, 06123 Perugia, ITALY

E-mail: matevin@ipguniv.bitnet

BEDNAREK Roman , Institute of Mathematics, A. Mickiewicz University, Faculty of Mathematics and Computer Science, 60-769 Pozna´, POLAND

E-mail: bednarek@plpuam11.bitnet

BELLONI Marino, Dipartimento di Matematica, Universitá di Pisa, Via F. Buonarroti 2, 56127 Pisa, ITALY

E-mail: belloni@gauss.dm.unipi.it

BERGER Günter, Fakultät für Mathematik und Informatik, Mathematisches Institut, Abteilung Analysis/Geometrie, Augustusplatz 10-11, D-04109 Leipzig, GERMANY

BOGNÁR Gabriella, Mathematical Institute, Mechanical University of Heavy Industries, Miskolc University, H-3515 Miskolc-Egyetemváros, HUNGARY

E-mail: h5354bog@ella.hu

BRIGOLA Rolf, Fachhochschule Nürnberg, Bereich Allgemeinwissenschaften und Informatik, Kesslerplatz 12, D-90121 Nürnberg, GERMANY

E-mail: brigola@cd4680fs.rrze.uni-erlangen.de

BROCK Friedemann, Universität Leipzig, Institut für Mathematik, Augustusplatz 9-10, D-04109 Leipzig, GERMANY

E-mail: brock@mathematik.uni-leipzig.d400.de

BROWN Richard C., Department of Mathematics, University of Alabama, Box 870350, Tuscaloosa, AL 35487-0350, U.S.A.

E-mail: dicbrown@ua1vm.ua.edu

BUGAJEWSKI Dariusz, Faculty of Mathematics and Informatics, A. Mickiewicz University, Matejki 48/49, 60-769 Pozna´, POLAND

BURENKOV Viktor I., Department of Differential Equations, University of Friendship of People, Miklukho-Maklaya 6, Moscow 117198, RUSSIA

E-mail: satel@log.mian.su

CHIAPPINELLI Raffele, Dipartimento di Matematica, Universita della Calabria, 87036 Arcavacata di Rende (Cs), ITALY

E-mail: 2157chi@icsuniv

CHIARENZA Filippo, Dipartimento di Matematica, Città Universitaria, V.le A. Doria 6, 95125 Catania, ITALY

E-mail: chiarenza@dipmat.unict.it

CIANCHI Andrea , Instituto di Matematica, Facoltà di Architettura Università di Firenze, Via dell’Agnolo 14, 50122 Firenze, ITALIA

CICOGNANI Massimo, Dipartimento di Matematica, Università di Ferrara, Via Machiavelli 35, 44100 Ferrara, ITALY

E-mail: cic@ifeuniv.unife.it

DANĚČEK Josef, Institute of Mathematics, FAST VUT, Žižkova 17, 602 00 Brno, CZECH REPUBLIC

DRÁBEK Pavel, Department of Mathematics, University of West Bohemia, Faculty of Applied Sciences, P.O. Box 314, 306 14 Plzeň, CZECH REPUBLIC

E-mail: pdrabek@.zcu.cz

EDMUNDS David E., Centre for Mathematical Analysis, University of Sussex, Falmer, Brighton BN19QH, UNITED KINGDOM

E-mail: D.E.Edmunds@sussex.ac.uk

FARKAS Walter, Johannes Gutenberg Universität, Fachbereich 17, Mathematik, Postfach 3980, D-55099 Mainz, GERMANY (current address)

E-mail: farkas@mat.mathematik.uni-mainz.de
       farkas@roimar.imar.ro

FERONE Adele, Università di Napoli “Federico II”, Dipartimento di Matematica e Applicazzione, Compleso Monte Sant Angelo Edificio ”T”, Via Cintia, 80126 Napoli, ITALY

E-mail: eferone@matna2.dma.unina.it

FOURNIER Richard, Department of Mathematics, Dawson College, 3040 Sherbrooke Street West, Montréal (QC) H3Z 1A4, CANADA

E-mail: fournric@ere.umontreal.ca

FRANCŮ Jan, Department of Mathematics, FME, Technical University, Technická 2, 616 69 Brno, CZECH REPUBLIC

E-mail: francu@kinf.fme.vutbr.cz

FREDDI Lorenzo, Dipartimento di Matematica e Informatica, Via Zanon 6, 33100 Udine, ITALY

E-mail: freddi@dm.unipi.it

FUJII Nobuhiko, Department of Mathematics, Tokai University (Shimizu) 3-20-1 Orido, Shimizu, Shizuoka 424, JAPAN

E-mail: nfujii@simizugw.cc.u-tokai.ac.jp

GALAZ-FONTES Fernando, CIMAT, Apartado Postal 402, 36 000 Guanajuato Gto., MÉXICO

E-mail: cimat@redvax1.dgsca.unam.mx

GARCÍA-CUERVA José, Departamento de Matemáticas C-XVI, Facultad de Ciencias, Universidad Autónoma de Madrid, 28049 Madrid, SPAIN

E-mail: cuerva@vm1.sdi.uam.es

GARCÍA DEL AMO Alejandro, Departamento de Analisis Matematico, Facultad de Ciencias Matematicas, Universidad Complutense de Madrid, 28040 Madrid, SPAIN

E-mail: garciada@mat.ucm.es

GOGATISHVILI Amiran, Institute of Mathematics, Georgian Academy of Sciences, Z. Ruchadze 1, Tbilisi 380093, REPUBLIC OF GEORGIA

E-mail: gogati@imath.kheta.georgia.su

GOL’DMAN Mikhail L., Krupskaya Str. 8-1-187, 117311 Moscow, RUS-SIA

E-mail: seulydia@glas.acp.org

GURKA Petr, Department of Mathematics, Agricultural University, 160 21 Prague 6, CZECH REPUBLIC

E-mail: gurka@tech.vsz.cz

GUTIERREZ Cristian E., Department of Mathematics, Temple University 038-16, Philadelphia, PA 19122, U.S.A.

E-mail: gutier@euclid.math.temple.edu

HARLÉN Mikael, Chalmers Intitute of Technology 41296 Göteborg, SWE-DEN

E-mail: harlen@math.chalmers.se

HAROSKE Dorothee, Fakultät für Mathematik und Informatik, FSU Jena, UHH, 17 OG, D-07740 Jena, GERMANY

E-mail: haroske@minet.uni-jena.de

HARRIS Desmond J., School of Mathematics, University of Wales, College of Cardiff, Senghennydd Road, P.O.Box 926, Cardiff CF2 4YH, UNITED KINGDOM

ILYIN Alexei, Hydrometeorological Research Center, Predtechentskyi 9–13, Moscow 123 376, RUSSIA

JOHN Oldřich, Faculty of Mathematics and Physics, Charles University, Sokolovská 86, 186 00 Praha 8, CZECH REPUBLIC

E-mail: john@karlin.mff.cuni.cz

KALITA Evgenii A., Institute of Applied Mathematics and Mechanics, Academy of Sciences of Ukraine, Rosa Luxemburg str. 74, Donetsk 340 114, UKRAINE

KAWOHL Bernd, Mathematisches Institut, Universität zu Köln, D-50923 Köln, GERMANY

E-mail: kawohl@mi.uni-koeln.de

KITTILÄ Arto, Department of Mathematics, University of Oulu, Linnanmaa, SF-90570 Oulu, FINLAND

E-mail: kittila@cc.oulu.fi

KOKILASHVILI Vakhtang, Institute of Mathematics, Georgian Academy of Sciences, Z. Ruchadze 1, Tbilisi 380093, REPUBLIC OF GEORGIA

E-mail: kokil@imath.georgia.su

KONEČNÝ Milan, 739 44 Brušperk 500, CZECH REPUBLIC

E-mail: konecny@oudec.osu.cz

KOSKELA Pekka, Department of Mathematics, University of Jyväskylä, P.O. Box 35, SF-40351 Jyväskylä, FINLAND

E-mail: pkoskela@finyu

KOSTENKO Irina S., Department of Numerical Mathematics, Lvov Polytechnical Institute, Stepan Bandera str.1, Lvov 290 000 UKRAINE KOVÁČIK Ondrej, Department of Mathematics, University of Transport and Communications, Faculty of Civil Engineering, J.M. Hurbana 15, 010 26 Žilina, SLOVAK REPUBLIC

KOWALEWSKI Wojciech, Institute of Mathematics, A. Mickiewicz University, Faculty of Mathematics and Computer Science, Matejki 48/49, 60-769 Pozna´, POLAND

E-mail: fraktal@plpuam11.bitnet

KRBEC Miroslav, Mathematical Institute, Czech Acad. Sci., Žitná 25, 155 67 Prague 1, CZECH REPUBLIC

E-mail: krbecm@earn.cvut.cz

KUFNER Alois, Mathematical Institute, Czech Acad. Sci., Žitná 25, 155 67 Prague 1, CZECH REPUBLIC

E-mail: kufner@earn.cvut.cz

LANG Jan, Mathematical Institute, Czech Acad. Sci., Žitná 25, 155 67 Prague 1, CZECH REPUBLIC

E-mail: lang@mff.karlin.cz

LEINFELDER Herbert, Fachhochschule Nürnberg, Bereich Allgemeinwissenschaften und Informatik, Kesslerplatz 12, D-90121 Nürnberg, GERMANY

E-mail: leinfelder@cd4680fs.rrze.uni-erlangen.de

LORENTE María, Análisis Matemático, Faculdad de Ciaencias, Universidad de Málaga, Campus de Teatinos, 29071 Málaga SPAIN

E-mail: m lorente@ccuma.uma.es

LOULIT Ahmed, Université Libre de Bruxelles, Départament de Mathématique CP 214, Campus de la Plaine, 1050 Bruxelles, Belgium

E-mail: ulbmath@ulb.ac.be

LOVE Eric Russell, 63 Yarrbat Avenue, Balwyn, Victoria 3103, AUSTRALIA

MALIGRANDA Lech, University of Luleå, Department of Mathematics, S-97187 Luleå, SWEDEN

E-mail: lech@sm.luth.se

MARTÍN-REYES Francisco J., Analysis Matemático, Facultad de Ciencas, Universidad de Málaga, Campus de Teatinos, 29071 Málaga, SPAIN

E-mail: martin reyes@ccuma.uma.es

McINTOSH Alan, School of Mathematics and Physics, Macquarie University, North Ryde, Sydney, N.S.W. 2109, AUSTRALIA

E-mail: alan@macadam.mpce.mq.edu.au

MÍKA Stanislav, Department of Mathematics, University of West Bohemia, Faculty of Applied Sciences, P.O. Box 314, 306 14 Plzeň, CZECH REPUBLIC

E-mail: mika@czu.cz

MOHAPATRA Ram N., Department of Mathematics, University of Central Florida, Orlando, Florida 32816, U.S.A.

MOŠOVÁ Vratislava, Na vyhlídce 14, 682 01 Vyškov, CZECH REPUB-LIC

NAKAMURA Akihiro, Tokai University, School of High Technology for Human Welfare, 317 Nishino Numazu Shozuoka 410–03, JAPAN

NEKVINDA Aleš, Department of Mathematics, Faculty of Civil Engineering, Prague Technical University, Thákurova 7, 166 29 Praha 6, CZECH REPUBLIC

NOWAK Marian, Faculty of Mathematics and Informatics, Adam Mickiewicz University, Matejki 48/49, 60-769 Pozna´, POLAND

E-mail: nowakmar@plpuam11.bitnet

OINAROV Ryskul, Institute of Mathematics and Mechanics, Kazakh Academy of Sciences, Pushkin 125, Alma-Ata 480021, REPUBLIC OF KAZAKHSTAN

E-mail: oinarov@lma.itpm.alma-ata.su

OPIC Bohumír, Mathematical Institute, Czech Acad. Sci., Žitná 25, 155 67 Prague 1, CZECH REPUBLIC

E-mail: opic@earn.cvut.cz

PÉREZ Carlos, Departamento de Matemáticas, Universidad Autónoma de Madrid, 28049 Madrid, SPAIN

E-mail: cperezmo@ccuam3.sdi.uam.es

PERSSON Lars-Erik, University of Luleå, Department of Mathematics, S-95187 Luleå, SWEDEN

E-mail: larserik@sm.luth.se

PICK Luboš, Mathematical Institute, Czech Acad. Sci., Žitná 25, 155 67 Prague 1, CZECH REPUBLIC

E-mail: pick@earn.cvut.cz

POSTERARO M. Rosaria, Dipartimento di Matematica e Applicazzione “R. Caccioppoli”, Complesso Monte S. Angelo, Via Cintia, 80126 Napoli, ITALY

E-mail: posterar@matna1.dma.unina.it

RÁKOSNÍK Jiří, Mathematical Institute, Czech Acad. Sci., Žitná 25, 155 67 Prague 1, CZECH REPUBLIC

E-mail: rakosnik@earn.cvut.cz

SARAIVA Luis, CMAF, Av. Prof. Gama Pinto, 2, 1699 Lisboa Cedex, PORTUGAL

E-mail: mmff5@ptmat.lmc.fc.ul.pt

SAWYER Eric T. , Department of Mathematics and Statistics, McMaster University, Hamilton, Ontario L8S 4K1, CANADA

E-mail: sawyer@mcmail.cis.mcmaster.ca

SCHMEISSER Hans-Jürgen, Fakultät für Mathematik und Informatik, FSU Jena, UHH, 17 OG, D-07740 Jena, GERMANY

E-mail: schmeiszer@minet.uni-jena.de

SCHOTT Thomas, Fakultät für Mathematik und Informatik, FSU Jena, UHH, 17 OG, D-07740 Jena, GERMANY

E-mail: schott@minet.uni-jena.de

SICKEL Winfried, Fakultät für Mathematik und Informatik, FSU Jena, UHH, 17 OG, D-07740 Jena, GERMANY

E-mail: sickel@minet.uni-jena.de

SKRZYPCZAK Leszek, Institute of Mathematics, A. Mickiewicz University, Faculty of Mathematics and Computer Science, Matejki 48/49, 60-769 Pozna´, POLAND

E-mail: lskrzyp@plpuam11.bitnet

STANOYEVITCH Alexander, Department of Mathematics, University of Hawaii, Honolulu 96822–2273, U.S.A.

E-mail: alex@.math.hawaii.edu

STARÁ Jana, Faculty of Mathematics and Physics, Charles University, Sokolovská 86, 186 00 Praha 8, CZECH REPUBLIC

E-mail: stara@karlin.mff.cuni.cz

STEPANOV Vladimir D., Russisan Academy of Sciences, Far East Branch, Computing Center, Kim-Ju-Czen 65, 680 063 Khabarovsk 63, RUSSIA

E-mail: 1600@ac.khabarovsk.su

SUN Jim Quile, Department of Mathematics, The University of Western Ontario, Middlesex College, London, Ontario N6A 5B7, CANADA

E-mail: gsun@uwovax.uwo.ca

TADDIA Nicola, Dipartimento di Matematica, Università di Ferrara, Via Machiavelli 35, 44100 Ferrara, ITALY

E-mail: tad@ifeuniv.unife.it

TALENTI Giorgio, Dipartimento di Matematica “U. Dini” Università degli Studi di Firenze, Viale Morgagni 67/A, 50134 Firenze, ITALY

E-mail: talenti@udinimath.unifi.it

de la TORRE Alberto, Departamento de Analisis Matemáticas, Facultad de Ciencias, Universidad de Málaga, Campus de Teatinos, 29071 Málaga, SPAIN

E-mail: torre r@ccuma.uma.es

TUMASHOVA O.V., Department of Numerical Mathematics, Lvov Polytechnical Institute, Stepan Bandera str.1, Lvov 290 000, UKRAINE

VERDIEV Verdi, Engels Str. 52–10, Makhatszkala, RUSSIA

VINTI Gianluca, Dipartimento di Matematica, Università degli Studi di Perugia, Via Vanvitelli 7, 06123 Perugia, ITALY

E-mail: matevin@ipguniv.unipg.it

VOLPICELLI Roberta, Università di Napoli “Federico II”, Dipartimento di Matematica e Applicazzione, Compleso Monte Sant Angelo Edificio ”T”, Via Cintia, 80126 Napoli, ITALY

E-mail: volpicel@matna2.dma.unina.it

WHEEDEN Richard L., Department of Mathematics, Rutgers University, New Brunswick, NJ 08903, U.S.A.

E-mail: wheeden@math.rutgers.edu

ZACHARIAS Klaus, Institut für Angewandte Analysis und Stochastik (IAAS), Mohrenstr. 39, D-10117 Berlin, GERMANY

E-mail: zacharias@iaas-berlin.d400.de

ZANGHIRATI Luisa, Dipartimento di Matematica, Università degli Studi di Ferrara, Via Machiavelli 35, Ferrara, ITALY

E-mail: zan@ifeuniv.bitnet

ZIMA Mirosława, Department of Mathematics, Pedagogical University of Rzeszów, Rejtana 16A, 35–310 Rzeszów, POLAND